OpenStudy (anonymous):

how do you solve this equation ? y=2x-1 2x+2y=22

OpenStudy (anonymous):

y-2x=-1 2x+2y=22 _______________ 3y=21 y=7 _________ 2x+2(7)=22 2x+14=22 2x=22-14 2x=8 x=4

OpenStudy (anonymous):

is that a one solution, infently many solutions

OpenStudy (anonymous):

no this equation has one solution not infently many solutions .the equation will have infently many solutions when virables more than equation